A train moves in a speed of 300 K/H. If the speed is increased by 5 K/H the time will decrease by 2 hours. What is the speed of the train. Show your work.
Answer by satyareddy22(84)

speed(S)=300 k/h,Time(T)=t
when S is increased by 5 k/h then S = 300+5=305 k/h,
Time is decreased by 2 hours =>T =t-2
Distance = S * T
Distance is same in both cases
so, 300 t = 305(t-2)
300t = 305t - 305 * 2
305t - 300t = 610
5t=610
t=610/5 =122h